You need 20 liters of 20% acid solution. You have containers of 10% solution and 25% solution. How many liters of each should you combine to get the needed solution?
Answer by fractalier(6550)

Call the amount of 10% you use, x.
Then the amount of 25% solution will be 20-x.
The set up looks like this
x(.10) + (20-x)(.25) = 20(.20)
.1x + 5 - .25x = 4
-.15x = -1
x = 6.67 liters of 10% solution
20-x = 13.33 liters of 25% solution